OpenStudy (anonymous):

PLEASE HELP... Which of the following demonstrates the Commutative Property of addition? A. 4+2=2+4 B. 0(3+1)=0 C. (1+2)+3=1+(2+3) D. ax1=a

OpenStudy (anonymous):

a+b=b+a

OpenStudy (anonymous):

which option has that framework?

OpenStudy (anonymous):

There is no option for that answer.

OpenStudy (anonymous):

no, that's just the basic way to represent commutative addition

OpenStudy (anonymous):

an example would be if a=1 and b=2: 1+2=2+1

OpenStudy (anonymous):

now which answer choice is similar to the example?

OpenStudy (anonymous):

D would be, right?

OpenStudy (dtan5457):

D is not right, that is a identity property.

OpenStudy (dtan5457):

Commutative in shorter words is that, the order when adding does NOT matter. 1+2+3=6 2+3+1=6

OpenStudy (anonymous):

A. 4+2=2+4

OpenStudy (anonymous):

A is the answer?

OpenStudy (dtan5457):

Aww come on now, there was a reason we were explaining everything.

OpenStudy (anonymous):

The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around.